OpenStudy (anonymous):

how do i solve: 7x+4= sq.rt. of 96??

OpenStudy (anonymous):

7x+4=sqrt(96) => x=(sqrt(96)-4)/7

OpenStudy (radar):

\[7x +4=\sqrt{96}\]

OpenStudy (radar):

\[7x =\sqrt{96}-4\]

OpenStudy (radar):

\[7x=\sqrt{16\times6}-4\]

OpenStudy (radar):

\[7x=4\sqrt{6}-4\] then \[x=(4\sqrt{6}-4)/7\]